OpenStudy (anonymous):

solve cos(2beta-pi/2)=-1 where 0

OpenStudy (ranga):

For what angle is cosine -1?

OpenStudy (anonymous):

cos(pi)=-1

OpenStudy (ranga):

correct. Equate 2beta - pi/2 and pi and solve for beta.

OpenStudy (anonymous):

cos(2beat-pi/2)=pi + 2pi k, where k is an integer

OpenStudy (anonymous):

how do I get cos beta by it self?

OpenStudy (ranga):

Yes. Solve for beta

OpenStudy (ranga):

cos(2beta-pi/2) = cos(pi + 2pi k), where k is an integer. Therefore, (2beta-pi/2) = (pi + 2pi k) Solve for beta.

OpenStudy (anonymous):

oh i got it thanks!!!

OpenStudy (ranga):

If cos(A) = cos(B) then A = B + 2kpi where k is any integer (positive, zero or negative). Here, cos(2beta - pi/2) = cos(pi) Therefore, 2beta - pi/2 = pi + 2kpi. Add pi/2 to both sides: 2beta = pi + 2kpi + pi/2 = 3/2pi + 2kpi divide both sides by 2: beta = 3/4pi + kpi Here we want beta to be in the range [0, 2pi]. Negative integer values for k will yield negative angle for beta and therefore can be ignored. Try k = 0, 1, 2 and stop when it goes past 2pi. You are welcome.